OpenStudy (anonymous):

integral of x/(25 - x^2)^1/2

OpenStudy (anonymous):

\[\int\limits_{?}^{?} x/\sqrt{25 - x^{2}}\]

OpenStudy (ybarrap):

using substitution: -sqrt(25-x^2) + C, C a constant

OpenStudy (anonymous):

use trig sub like last time = you have the same thing only backwards

OpenStudy (anonymous):

use the substitution x=asin(x)

OpenStudy (anonymous):

could i do x = 5cos(theta)? then dx = -5sin(theta)?

OpenStudy (anonymous):

x=5sin(theta)

OpenStudy (anonymous):

there is a special trig relationship i'm not sure what it is you'd have to ask jim thompson or someone who really knows his stuff. it carries throughout calc as you can see it in inverses like i said before

OpenStudy (ybarrap):

let u = 25 - x^2 then du = -2x so integral -1/2 du/u^(1/2) = - u^(1/2) substituting back: Ans -sqrt(25-x^2) + constant

OpenStudy (anonymous):

give me a sec i'll attach a pic

OpenStudy (anonymous):